Find the integral surface of the equation:


(x²-yz)p+(y²-zx)q =z²-xy


passing through the line x=1 , y=0

1
Expert's answer
2022-01-27T00:46:57-0500

"\text{To find the integral surface of the given equation, we proceed as follows:}"

The auxiliary equations are;

"\displaystyle
\frac{dx}{x^2-yz}=\frac{dy}{y^2-zx}=\frac{dz}{z^2-xy}\\"

"\displaystyle
\Rightarrow \frac{dx-dy}{x^2-yz-y^2+zx}=\frac{dy-dz}{y^2-zx-z^2+xy}=\frac{dz-dx}{z^2-xy-x^2+yz}\\
\Rightarrow \frac{dx-dy}{(x-y)(x+y+z)}=\frac{dy-dz}{(x+y+z)(y-z)}=\frac{dz-dx}{(x+y+z)(z-x)}\\
\Rightarrow \frac{dx-dy}{x-y}=\frac{dy-dz}{y-z}=\frac{dz-dx}{z-x}\cdots\cdots\cdots\cdots(1)\\"

Integrating first two members of (1) yields;

"\displaystyle
\log(x-y)=\log(y-z)+\log c_1\\
\Rightarrow\log \frac{x-y}{y-z}=\log c_1\\
\Rightarrow \frac{x-y}{y-z}=c_1\\"

Similarly from the last two members of (1), we have;

"\displaystyle
\frac{y-z}{z-x}=c_2"

The required solution is;

"\displaystyle
f\left[\frac{x-y}{y-z},\frac{y-z}{z-x}\right]=0"